I need help solving this problem using cramer's rule
-3x-4y=9
9x+10y=-3

Answer by checkley77(12844)

-3x-4y=9
9x+10y=-3
******************
9/-3 -4/10
----------- = x
-3/9 -4/10
********************
9*10-(-3)(-4)
--------------- = x
-3(10)-(9*-4)
**************** =78/6=13 ans. for x.
(90-12)/(-30+36)
*********************
-3/9 9/-3
---------- = y
-3/9 -4/10
********************
-3*-3-9*9
----------- = 9-81/-30+36=-72/6=-12 ans for y
-3*10-9*-4
***********
Proof:
-3*-13-4*12=9
-39+48=9
9=9
9*13+10*-12=-3
117-120=-3
-3=-3